OpenStudy (anonymous):

Which equation can be used to find 27% of 255? 27/100 = x/225 27/225= x/100 100/225 =x/27 x/27=100/225

OpenStudy (anonymous):

\( \Large x=\frac{27}{100}\times225 \) \(\Large \Rightarrow \frac{x}{225}=\frac{27}{100} \)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

:) Welcome